### v1

@@ Abstract
We explore the enhancement of Human-in-the-Loop video annotation by integrating automatic capabilities to ease the task for annotators and assess their performance. The research delves into the practical implications of the annotation processes, the integration of AI components, and the evaluation of its outcomes. We analyze their impact on efficiency, accuracy, and overall annotation quality.
Focusing on the Human-in-the-Loop for video annotation tasks, we implemented a single-iteration scheme using Label Studio and AI-powered zero-shot pre-annotations. Using this framework, we designed a test based on the annotation of the UCF-Crime dataset to discriminate between normal and abnormal activities in video footage.
Our results evidence how automatic AI-based pre-annotation can streamline the video annotation workflow, empowering human annotators and optimizing the overall pipeline. Using the pre-annotated data, we observed a 35% reduction in the annotation time for 70% of the annotators with similar quality annotations, compared to the traditional manual annotation task. Results are consistent with asset duration and complexity. We also observed that while annotators rapidly learned to use the tool, the produced annotations are more coherent among annotators and better match the natural clustering of the video frames.  [^1] The collected data is accessible on https://github.com/jgnav/AI-Boosted-Video-Annotation/.
@@ 1 Introduction
Video annotation has been the cornerstone of data-driven learning, particularly in scenarios demanding highly reliable annotated data. Alleviating the burden for annotators, expediting the process, and improving the annotations are recurrent goals to which we want to contribute.
By incorporating Artificial Intelligence (AI) tools into the annotation process within the Human-in-the-Loop (HITL) paradigm, the aim is to reduce the number of frames to annotate and make each human input more valuable, reducing the work time and minimizing the probability of an error or inconsistency. We are leveraging human intelligence to handle the nuances and complex decisions, which are the weak points of fully automated annotation, and harnessing the machines’ ability to process large amounts of data.
We devised a comprehensive annotation test to assess pre-annotations’ impact on the video annotation flow. We aim to explore the advantages and disadvantages of this approach and its impact on final data outcomes. We have decided to conduct the test using real-world videos primarily captured from surveillance cameras. Surveillance videos are characterized by their typically low resolution and poor quality, recorded from a stationary point amidst varying lighting and environmental conditions. This choice is because it presents a challenging use case for pre-annotating automatically with AI models and is a widely valued scenario in today’s context.
Additionally, we chose a group of annotators for data labeling instead of relying on a limited set of “experts”. This approach leverages the collective intelligence of a diverse pool of individuals for tasks such as labeling, classification, and data annotation. Nonetheless, it would be interesting to evaluate the results of a similar test on professional annotators to compute performance figures and evaluate their proficiency level. For the time being, this was out of the scope of our investigation.
Crowdsourcing annotations for discrete labeling has become a major trend these days. The task involves collecting multiple judgments from annotators, which are then aggregated to establish a single “ground truth” published in the dataset [6]. However, it is crucial to recognize that disagreements among annotators may contain valuable nuances worth exploring. Our test follow a crowdsourcing methodology, implementing stringent quality control measures to ensure the generation of high-quality annotations. These measures include mandatory training on the annotation tool before the test, removing outliers (annotations that deviate logically from the expected average), and assigning each video for annotation with a 6-vote overlap. This ensures that each task receives input from six different annotators, promoting diversity in perspectives.
Therefore, our primary objective is to design a hypothetical annotation test that affords annotators significant freedom, allowing for the consideration of diverse criteria. Subsequently, we conducted a study on and analyzed the impact of our pre-annotations on annotators’ criteria and efficiency and evaluated the consistency of the results obtained.
@@ 2 Related work
The literature describes several different annotation tasks and platforms. Reviewing previous initiatives incorporating AI-based pre-annotations to streamline the annotation process, we observe the inherent complexity and diversity, adding a layer of intricacy to the evaluation. However, let us first clarify the concept of data annotation: Data annotation involves the labeling or tagging data with specific labels or tags, rendering it more comprehensible for machines. This crucial activity is pivotal in developing supervised Machine Learning (ML) algorithms and models and validating their unsupervised and self-supervised counterparts [26].
These annotations vary significantly in terms of the task at hand and the tools used, and these aspects must be considered in the annotation process.
2.1 Common annotation tasks
Video assets represent rich sources of information, prevalent across a diverse range of applications and tasks associated with information management. The extensive variety of video annotation tasks contributes to a broad spectrum of annotation tasks.
The granularity required for video annotation gives rise to many annotation cases. The annotation may encompass the entire video [21, 27], temporal segments (shots) [27, 1], frames (temporal segments with zero duration) [9, 32], regions within frames [20, 19], or even moving regions—referring to a region tracked across a sequence of frames [4].
Another important use case relates to the Video Moment Retrieval (VMR). VMR retrieves a segment described by a given natural language query from an untrimmed video. This task, alternatively known as natural language video localization and video temporal grounding, is a fundamental challenge in comprehending computer vision and visual information retrieval. VMR finds wide applications in scenarios such as video browsing websites and semantics-based video search engines [30, 3, 2]. Interestingly, the annotations introduced for each use case vary significantly, and therefore, we must consider the task while looking for AI tools that may boost the annotation process.
2.2 Tools for video annotation
We have reviewed the characteristics of the primary video annotation platforms. CVAT, an open-source annotation platform developed by Intel, offers a comprehensive set of annotation tools, including object detection, image segmentation, and video annotation. CVAT supports multiple annotation formats and provides APIs for easy integration with other computer vision applications and workflows. Being a web-based tool, CVAT can be downloaded and installed on custom local or cloud servers, making it customizable and extensible. It also includes tools for medical image conversion and annotation [16].
ANU CVML VIDeo Annotation Tool (VIDAT) is a high-quality, simple, and efficient open-source video annotation tool developed by ANU Computer Vision Machine Learning. It is user-friendly, supporting multiple annotation types, including temporal segments, object bounding boxes, semantic and instance regions, tracklets, and human pose (skeleton). VIDAT runs in a browser without external libraries or the need for server-side processing. However, it is easy to integrate with a back-end “Human-in-the-Loop” processing. It ensures data security as the data does not need to leave the local machine, and it is open-source [14].
The Video and Image Annotation (VIA) is a lightweight, standalone, and offline software package that operates in a web browser without requiring installation or setup. Users can annotate at various levels of granularity, from the entire video down to specific segments, moving regions, individual frames, and even static regions within a frame. VIA also supports collaborative annotation of a large dataset by a group of human annotators. The BSD open-source license allows its use in academic projects or commercial applications [7].
The IBM VideoAnnEx annotation tool focuses on video annotation utilizing MPEG-7 metadata. Despite the tool no longer being actively maintained, VideoAnnEx remains accessible and serves as an illustrative example of content annotation aligned with the MPEG-7 initiative. While its maintenance status may pose considerations for long-term use, the tool offers valuable insights and a historical perspective on content annotation methodologies within the MPEG-7 framework [15].
Finally, despite the features described, our choice for the tests was Label Studio for several compelling reasons. Firstly, Label Studio has proven to be a reliable platform with a track record of annotating high-quality data. It is free and open-source and allows extensive customization options for the annotation interface. Moreover, it provides essential functionalities for our test, such as frame classification over a temporal bar and the automatic, precise measurement of annotation time per video [24].
2.3 AI-boosted annotation
Video annotation constitutes a time-consuming and labor-intensive endeavor, prompting the exploration of various approaches to enhance efficiency through integrating the Human-in-the-Loop paradigm. The primary goal is to reduce the number of frames requiring annotation, thus maximizing the value of each human input in the process. The expansive landscape of video annotation applications has given rise to diverse methodologies tailored to specific use cases.
In [13], a unified framework for generic video annotation with bounding boxes is introduced. This model combines interpolating and extrapolating capabilities, tracking target appearance in multiple annotated frames. Additionally, it proposes a guiding mechanism that sequentially generates suggestions for the next frame based on previous annotations.
The work presented in [3] focuses on video moment retrieval, aiming to identify the start and end timestamps of a moment in a video based on a natural language query. The introduced paradigm, “glance annotation”, requires only the timestamp of a single random frame, referred to as a “glance”, within the temporal boundary of the fully supervised counterpart. ViGA, utilizing contrastive learning, divides the input video into clips and contrasts them with queries, assigning glance-guided Gaussian distributed weights to all clips.
In [5], a data analysis framework is proposed to facilitate the interactive annotation of racket sports videos with the support of computer vision algorithms. This framework leverages machine learning models in computer vision to help users extract essential events from videos and provides interactive tools for data annotation.
A lightweight approach named Video2Action is presented in [8]. In video creation, where annotating actions is a common requirement, Video2Action automatically generates action scenes and predicts action locations in the video using image processing and deep-learning methods.
Other authors have developed their AI-powered tools to ease video annotation. One of these is the pre-annotation tool developed by the authors of the present contribution using contrastive learning and cross-modal representations (to be published).
3 Annotation Assessment
The annotation process focused on classifying frames from surveillance camera videos containing anomalous events. This task requires annotators to categorize and group the contents of each video. It is important to acknowledge that the criteria for classification may vary among annotators, and there may not be a single valid criterion, even though we defined a single task.
Annotators interacted with a video interface that featured the video itself and a temporal bar, facilitating the convenient classification of frames into five distinct classes. The interface also includes checkboxes for binary selection on each class, allowing annotators to indicate whether the video segment corresponds to an abnormal event (see Figure 1).
Figure 1: Label Studio interface [24] as adapted for the performed experiences and test.
Our annotation pipeline generates pre-annotations, providing comprehensive annotations for each video. Additionally, keyframes representing each class serve as visual references during the annotation process. Therefore, there are two types of annotations: without pre-annotation or unassisted, and with pre-annotation. In the second case, the annotator is presented with the fully pre-annotated video, and their task is to review the pre-annotation and modify it as they deem fit according to their criteria. We generated these pre-annotations using multi-modal AI models incorporating both vision and text. This work is still pending publication.
Our objective is to study the differences in annotations with and without pre-annotation and to identify any potential biases introduced by pre-annotations. Another crucial aspect under consideration is annotation time; our pipeline aims to improve the annotation process’s efficiency significantly.
For a more in-depth description of the annotation process, our CrowdWorkSheets responses [6] have been included in Appendix 0.A, a framework widely employed in similar works [10].
3.1 Experiment setup
We selected Label Studio as our annotation platform for this study [24]. To carry out the annotation process, we engaged the participation of 18 voluntary annotators, all affiliated with a university. Each annotator was assigned ten different videos, with a unique combination of five videos containing pre-annotations and five without pre-annotations, presented in varying orders. Half of the annotators began with pre-annotated videos, while the other half started with videos without pre-annotations. The approximate time required per annotator was 30 minutes, excluding training. In total, 30 videos underwent annotation, each receiving six annotations from six different annotators (6-vote overlap) - three with pre-annotations and three without.
The video selection process involved a random sampling from all predefined annotation classes. Two filters were applied during the selection: the first filter was based on video duration, discarding those lasting over 40 seconds due to time constraints for annotators. We aimed for a diverse set of videos with shorter durations. The second filter excluded videos with explicit content, determined by the consensus of three researchers.
The entire annotation process was overseen by Sigma Cognition S.L. [11], an entity with extensive experience in annotation tasks. Their supervision ensured the quality and consistency of the annotations throughout the project.
3.2 Dataset of Real-world Videos
The dataset chosen for this task is UCF-Crime, which is recognized as one of the most comprehensive and diverse collections of surveillance videos. This dataset has been extensively utilized in research on this specific type of video domain [28, 12, 31].
The dataset comprises a total of 128 hours of videos, consisting of 1900 long untrimmed real-world surveillance videos. It covers 13 realistic anomalies, including fights, traffic accidents, burglaries, robberies, and more. These anomalies are significant due to their impact on public safety [22].
4 Results
An extensive analysis of the annotation test results has been conducted and divided into three main blocks. Figure 2 illustrates the framework of all the result evaluation processes carried out, including the inputs and outputs of each. Two primary evaluation metrics are extracted from the annotation process: frame classification and the annotation time per video for each annotator. Concerning the temporal analysis, the objective has been to measure if there has been a substantial gain in annotation time.
Within the quality evaluation block, an assessment was made to determine whether there was a noticeable change in the quality of annotations, examining their consistency and the potential introduction of biases. Finally, in the evaluation block utilizing embeddings representing the semantic information of the frames, the classification quality is assessed about the video’s semantic content.
Pre-anno-tationsAnnotationProcessTimestampsTemporalEvaluationTemporalAnalysisQualityAnalysisVisualSimilarityAggregation&HomogeneitySemanticLatent SpaceOperationsLabeled ClassesFrames EmbeddingsCLIPImageEncoderAI-BasedPre-annotationPipelineVideos
Figure 2: Diagram of the process followed to evaluate the results of the annotation test and to run the experiences. Videos (to the left) are automatically pre-annotated. These pre-annotations help the annotation process, while input frames are processed using the CLIP image transformer [18]. Annotation results (timestamps, labeled classes) and frame embeddings are used to assess annotation results.
4.1 Temporal analysis of annotations
The most crucial metric in this test is the temporal one, as the primary goal is to optimize the annotation time. To measure this, the annotation platform Label Studio accurately tracks the time elapsed from when the annotator opens the task until they click the submit button.
Given the number of annotations obtained, it is crucial to consider specific factors that may influence the temporal metric, especially concerning the presence of human factors and video-related elements in the data. Nonetheless, our focus is on studying discernible trends. More issues that were unrelated to the test, such as technical problems, annotator distractions, or those arising from inexperience, have been addressed separately by identifying outliers and assessing whether it is worthwhile to eliminate them from the annotations.
Figure 3 illustrates the distributions of all annotation data. The kernels of these distributions exhibit a behaviour similar a Gamma functions. Table 1 provides a comprehensive overview of the statistical analysis results. Parametric tests were carried out between the two annotation sets to evaluate their independence, and all tests affirmed their independence. On average, the annotation time is lower by 23.11% for cases where pre-annotation is applied.

### v2

@@ Abstract
Manual annotation remains the gold standard for high-quality, dense temporal video datasets, yet it is inherently time-consuming. Vision-language models can aid human annotators and expedite this process. We report on the impact of automatic Pre-Annotations from a tuned encoder on a Human-in-the-Loop labeling workflow for video footage. Quantitative analysis in a study of a single-iteration test involving 18 volunteers demonstrates that our workflow reduced annotation time by 35% for the majority (72%) of the participants. Beyond efficiency, we provide a rigorous framework for benchmarking AI-assisted workflows that quantifies trade-offs between algorithmic speed and the integrity of human verification. Code and collected interaction data are available at this repository.
Human-in-the-Loop (HITL), Evaluation Methodology, Video Annotation, Cross-Modal Learning
@@ 1 Introduction
Progress in video understanding has been accelerated by large-scale pretraining and, more recently, by Vision-Language Models (VLMs) that align visual content with natural language supervision (Radford et al., 2021; Wang et al., 2022; Zhao et al., 2024). However, for many downstream tasks that require supervised learning methods, including anomaly detection, temporal action localization, and open-world event understanding, the bottleneck remains the same: dense temporal annotation.
Unlike static image labeling, video annotation requires defining start and end timestamps for events that unfold over time, often with unclear boundaries, occlusions, and multi-actor interactions. This makes labeling slow, expensive, and inherently variable between annotators (Uma et al., 2021; Davani et al., 2022). The problem is particularly severe in high-ambiguity video footage, where events are not clearly defined and can be regarded as highly subjective, context-dependent, and culturally mediated. At the same time, model development (and fair evaluation) depends on high-quality annotated datasets (Gu et al., 2018; Grauman et al., 2022) and benchmarking (Sultani et al., 2018).
To mitigate labeling cost, the annotation industry has widely adopted Human-in-the-Loop (HITL) workflows: a model proposes Pre-Annotations and humans correct them, shifting the annotator’s role from “create” to “verify and refine”. Popular annotation platforms and production tools largely support this paradigm (CVAT.ai Corporation, 2023; Tkachenko et al., 2020). However, despite its prevalence, the research literature lacks rigorous, controlled studies quantifying how AI suggestions affect (i) annotation efficiency, (ii) inter-annotator consistency, and (iii) the semantic integrity of the resulting labels, especially in domains where ambiguity is the norm.
Speed alone is an incomplete objective. AI assistance can reduce time while quietly inducing semantic drift. This entails a systematic shift in human labeling toward the model’s priors, potentially reducing the validity of the label even as agreement increases. Prior work has explored ways to reduce human interaction in visual annotation (e.g., verification-centric labeling and interaction-efficient interfaces) (Papadopoulos et al., 2016, 2017; Kuznetsova et al., 2021; Deng et al., 2021; Feng et al., 2023), but there is limited evidence on the human decision dynamics under modern VLM-style Pre-Annotation for dense temporal segmentation in open-world footage.
In this work, we present and evaluate a hybrid HITL workflow for dense temporal video segmentation in a deliberately challenging, high-ambiguity setting. We use a resource-efficient cross-modal pipeline, based on a fine-tuned CLIP-style encoder and clustering (Rasheed et al., 2023; Hornik et al., 2012)., to generate temporally localized “strong clusters” that serve as Pre-Annotations for humans. Our focus is not on proposing yet another video model, but on providing a reproducible benchmarking framework for AI-assisted annotation that captures both throughput and label integrity. We conducted a controlled counterbalanced A/B study with $18$ annotators labeling videos from open-world datasets, instrumenting the interface to capture rich interaction telemetry (time, edits, label changes, and cluster modifications). Our contributions are:
  • A controlled stress test of AI-assisted dense temporal annotation in high-ambiguity footage, designed to isolate the effect of Pre-Annotations on human behavior.
  • An end-to-end hybrid workflow (Pre-Annotation + UI) that is lightweight and reproducible, enabling future work to benchmark human–model interaction under comparable conditions.
  • A multi-axis evaluation protocol that goes beyond speed, measuring changes in agreement/consensus and latent-space semantic validity, to detect potential semantic drift.
  • A public release of code and anonymized interaction traces to support reproducibility and follow-up research on human-centric dataset construction.
@@ 2 Related Work
AI-assisted annotation and Human-in-the-Loop (HITL).
Reducing human labeling costs has long been a practical driver of progress in computer vision. Early work showed that replacing full manual drawing with lightweight verification protocols can substantially reduce the annotation effort (Papadopoulos et al., 2016, 2017). Iterative refinement loops further reinforce the corrective interaction between the model and the annotator (Adhikari and Huttunen, 2021). For video, interaction-efficient systems have explored guidance via interpolation and frame selection (Kuznetsova et al., 2021), single-frame supervision for moment retrieval (Cui et al., 2022), and domain-specific tools to minimize human effort (Deng et al., 2021; Feng et al., 2023; Shrestha et al., 2023). In parallel, production tooling has operationalized model-assisted labeling at scale (e.g., CVAT and Label Studio) (CVAT.ai Corporation, 2023; Tkachenko et al., 2020). Despite this broad adoption, controlled evidence quantifying how Pre-Annotations alter human temporal segmentation behavior—beyond throughput—remains comparatively limited, especially in high-ambiguity settings.
Vision–language pretraining for efficient video representations.
Vision-language models such as CLIP provide transferable encoders that can be adapted to video tasks with relatively modest computation (Radford et al., 2021). Large-scale video foundation models define strong upper bounds in representation quality (Wang et al., 2022; Zhao et al., 2024), but are expensive to use as annotation engines. A complementary line of work studies efficient CLIP-style video learners, including frame-level encoding with temporal pooling and targeted finetuning (Rasheed et al., 2023; Wang et al., 2023). Our Pre-Annotation engine follows this resource-conscious philosophy and adds an explicit hierarchical structuring stage using spherical $k$-means (Hornik et al., 2012) to produce multi-granularity segments. Recursive retrieval methods such as RAPTOR (Sarthi et al., 2024), together with hierarchical video parsing approaches (Cui et al., 2022), provide a principled mechanism that reflects the intrinsic semantic structure of video content.
Anomaly detection and ambiguity.
Real-world anomaly detection benchmarks, such as UCF-Crime, are inherently open-ended and temporally ambiguous (Sultani et al., 2018). This ambiguity is compounded by known biases in open-world footage and by models’ tendency to exploit background cues rather than event semantics (Liu et al., 2022). Recent work also explores adapting vision–language models to weakly supervised video anomaly detection (Wu et al., 2024). These characteristics make open-world video a particularly demanding test domain for evaluating whether model suggestions help humans without silently steering them toward model priors.
Quality measurement beyond speed: agreement, drift, and semantics.
Annotation tool evaluations often emphasize time-per-label or proxy accuracy metrics (Liao et al., 2021). However, efficiency can mask semantic drift: systematic shifts in human boundaries toward the model’s proposed structure. To capture integrity, a common axis is inter-annotator agreement; more recent perspectives argue that disagreement can be meaningful signal rather than noise, especially in subjective tasks (Uma et al., 2021; Davani et al., 2022). We therefore pair agreement/consensus measures with latent-space validity signals inspired by embedding-based evaluation ideas in adjacent areas (Unterthiner et al., 2019; Shahapure and Nicholas, 2020). Finally, transparency frameworks such as CrowdWorkSheets encourage reporting annotator demographics, recruitment, and context to interpret human-centric findings responsibly (Díaz et al., 2022; Gustafson et al., 2023).
3 User Study and Workflow Design
3.1 Study Protocol
We conducted a controlled user study involving 18 volunteer annotators affiliated with a university institution. All participants completed a mandatory training session based on a standardized annotation guide to ensure consistent behavior. The protocol defined a dual-objective task: users were required to segment all distinct content within the footage—grouping frames that shared the same action, event type, or scenario —and then classify each segment. To reduce ambiguity, the guide explicitly defined ‘abnormal events’ as any ‘criminal, accidental, or catastrophic’ incident, with all other content defaulting to ‘normal’.
To control for attentional bias, the protocol mandated that annotators view each video in its entirety before beginning segmentation. Participants were instructed that frame-perfect precision was not the primary objective, with a target task duration of 8-12 minutes per video to simulate realistic, labeling conditions.
For evaluation, we randomly sampled 30 videos from the datasets (Section 4.1) test sets. The study utilized a counter-balanced A/B design. Each participant annotated a unique batch of 10 videos: 5 in the Assisted Condition (with generated Pre-Annotations) and 5 in the Unassisted Condition (raw footage). The order of conditions was randomized to decouple the tool’s “learning curve” from the impact of the suggestions, half the cohort began with Condition A and the other half with Condition B. The average session duration was approximately 60 minutes.
Given the inherent subjectivity of defining anomaly boundaries in untrimmed video, we did not rely on a single ground truth. Instead, we employed a 6-vote overlap crowdsourcing strategy. Each video in the test subset was annotated by six independent participants (three using the assisted workflow, three unassisted). This aggregation allows us to construct a Consensus Ground Truth by identifying the temporal segments where the majority of annotators agreed, filtering out individual noise and capturing the “natural” semantic boundaries of the events.
For full transparency regarding annotator recruitment, compensation, and ethical considerations, we have included a complete CrowdWorkSheets response in Appendix B. This documentation framework is widely employed in similar human-centric AI studies to ensure reproducibility and ethical rigor (Gustafson et al., 2023).
To manage the annotation task and compute pre-annotations, we deployed a hybrid workflow that integrates an adapted version of an open-source annotation tool with a resource-efficient cross-modal encoding pipeline. We utilize a proven retrieval framework—characterized by fine-tuned encoders and hierarchical clustering —as a controlled variable. This ensures that pre-annotations are grounded in a consistent semantic structure, allowing us to isolate and quantify the specific impact of algorithmic suggestion on human efficiency. The human interface had to be adapted for this specific task.
3.2 Providing a Graphical User Interface
We integrated our Pre-Annotation pipeline with Label Studio, an open-source data labeling platform selected for its flexibility and robust feature set (Tkachenko et al., 2020). The interface was customized to support a dense temporal annotation task where users categorize video segments into distinct classes. To facilitate reproducibility, the exact Label Studio interface configuration and template code are released alongside the project repository.
As illustrated in Figure 1, the interface comprises two primary components: a video rendering module for content inspection and a temporal control layer. This layer features two synchronized timelines: a video navigation bar and an annotation bar where segments are actively manipulated. The interface also includes playback speed controls to facilitate rapid scanning. Additionally, the interface includes controls for binary classification, enabling annotators to explicitly flag segments containing specific events (e.g., abnormal vs. normal activities).
(a) The main annotation workspace featuring the video player and class taxonomy controls.
(b) Detail of the temporal timeline showing pre-populated segments and binary selection toggles.
Figure 1: The customized Label Studio interface adapted for the hybrid workflow. Users interact with generated Pre-Annotations directly on the temporal bar (b), refining the suggested segments.
In the assisted workflow, the Pre-Annotation pipeline populates this timeline with suggested segments—the “strong clusters” derived from the Pre-Annotation pipeline—before the user begins. The annotator’s task effectively shifts from creating annotations from scratch to a verification and refinement process. Users must explicitly duplicate the read-only predictions into their active workspace, where they can adjust boundaries and correct class labels. To expedite this process, the interface supports keyboard shortcuts (e.g., Ctrl+C/V) for replicating segments. To further support this decision-making, the system provides keyframes—closest frame to the cluster center—representing each class as visual references during the process.
3.3 Producing the Pre-Annotations
We employ a cross-modal retrieval architecture explicitly adapted for video downstream tasks (Radford et al., 2021) to generate high-quality suggestions for the HITL workflow. This pipeline serves as the generator of “strong clusters”—semantically coherent video segments that are presented to annotators for verification.
We explicitly leverage a frame-video dataset approach, as it allows for the robust maintenance of semantic information while incorporating spatiotemporal cues without prohibitive computational costs. This resource-efficient strategy has been validated in prior works (Rasheed et al., 2023; Ma et al., 2022b; Luo et al., 2021).
Encoding and Training Procedure
Given a set of video-text pairs $(V,P)$, we treat each video clip $V_{i}\in\mathbb{R}^{N\times H\times W\times C}$ as a sequence of $N$ frames. To capture temporal consistency without overhead, we employ a temporal pooling strategy. Each frame in $V_{i}$ is encoded to yield frame-level features $z_{i}\in\mathbb{R}^{N\times D}$. We then obtain a video-level embedding $v_{i}\in\mathbb{R}^{D}$ by applying average pooling over the $N$ frames, aggregating temporal data into a single fixed-size representation (Rasheed et al., 2023).
In parallel, the corresponding text description $p_{i}$ is processed by the text encoder to produce an embedding $t_{i}\in\mathbb{R}^{D}$ in the shared feature space. The alignment between modalities is driven by the cosine similarity $sim(v_{i},t_{i})$. We fine-tune the encoders using a contrastive approach, optimizing the symmetric Cross-Entropy (CE) loss. For a batch size of $N$, the loss for video-to-text is defined as:
$\mathcal{L}_{v2t}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\!\big(\operatorname{sim}(v_{i},t_{i})/\tau\big)}{\sum_{j=1}^{N}\exp\!\big(\operatorname{sim}(v_{i},t_{j})/\tau\big)}$ (1)
To ensure the model is robust for both retrieval directions, we apply a Multi-Task Learning (MTL) objective. The total loss $\mathcal{L}_{\text{total}}$ averages the losses for both video-to-text and text-to-video directions:
$\mathcal{L}_{\text{total}}=\frac{1}{2}\left(\mathcal{L}_{v2t}+\mathcal{L}_{t2v}\right).$ (2)
Figure 2 illustrates our proposed approach. This optimization ensures that the embedding space maximizes the similarity between correct video-text pairs while minimizing it for incorrect ones (Dong et al., 2021). By training on video clips paired with detailed textual descriptions, the model learns to effectively address zero-shot and open-class scenarios. We further analyze the effectiveness of this training in Appendix A.3.
[…]
Figure 2: Architecture of the Pre-Annotation pipeline. Video frames are processed by the image encoder to generate frame-level embeddings, which are pooled to produce a video-level embedding. These are aligned with text embeddings via contrastive learning, creating the semantic foundation for the hierarchical clustering.
Figure 3: Hierarchical clustering of video content. Frames (left) are recursively clustered into a tree structure (right) using spherical $k$-means. This structure allows the Pre-Annotation pipeline to propose segments at different levels of granularity, which users then verify or refine.
Hierarchical Content Modeling
To produce the actual Pre-Annotation segments, we model the video feature space as a hierarchical tree depicted in Figure 3. This structure reflects the inherent granularity of video content, from broad environments to specific actions.
We construct this tree by recursively clustering the feature arrays using spherical $k$-means (Hornik et al., 2012). This variant maximizes cosine similarity between data points and centroids, aligning with the $n$-dimensional hyperspherical nature of the CLIP latent space (Liang et al., 2022). The objective function is defined as:
$\max_{r_{ij},\,\boldsymbol{\mu}_{j}}\sum_{i=1}^{p}\sum_{j=1}^{k}r_{ij}\,\frac{\boldsymbol{\mu}_{j}\cdot\mathbf{x}_{i}}{\lVert\boldsymbol{\mu}_{j}\rVert\;\lVert\mathbf{x}_{i}\rVert}$ (3)
subject to $r_{ij}\in\{0,1\}$ and $\sum_{j=1}^{k}r_{ij}=1$ for all $i$, where $\boldsymbol{\mu}_{j}$ is the centroid of cluster $j$, $\mathbf{x}_{i}$ is the $i$-th data point, and $r_{ij}=1$ if and only if $\mathbf{x}_{i}$ is assigned to cluster $j$. The process begins with high-level clusters, and progressively refines them into fine-grained nodes until a stopping criterion is met. To ensure the segments represent meaningful semantic units rather than arbitrary temporal splits, we employ a maximum tree depth constraint (set to $T_{max}$=3) and a semantic cohesion threshold, halting the recursive subdivision when the cosine similarity within a cluster exceeds $\tau=$0.85. The resulting leaf nodes represent “strong clusters” that we interpret as distinct video segments. These segments constitute the Pre-Annotations provided to the human annotators and serve as the starting point for our evaluation.
4 Evaluation & Analysis
To provide a complete assessment of the hybrid workflow, we evaluate the results across three distinct dimensions: Efficiency (temporal gains), Consistency (inter-annotator agreement), and Standardization (structural coherence). Figure 4 outlines the evaluation framework, illustrating how raw logs, class labels, and frame embeddings are processed to derive these metrics.
Pre-anno-tationsAnnotationProcessTimestampsTemporalEvaluationTemporalAnalysisQualityAnalysisStructuralConsistencyAggregation&HomogeneitySemanticLatent SpaceOperationsLabeled ClassesFrames EmbeddingsImageEncoderPre-AnnotationPipelineVideos
Figure 4: Evaluation framework. The diagram maps the flow from the Pre-Annotation Pipeline (black blocks) to the three pillars of analysis: Temporal Efficiency, Homogeneity, and Structural Consistency.
4.1 Training Datasets & Setup
Datasets. We utilized a diverse composite of four datasets to ensure the encoder’s robustness across varying temporal structures and semantic contexts. UCF-Crime (Sultani et al., 2018) serves as the foundational domain, providing 1,900 untrimmed real-world surveillance videos (128 hours) characterized by low resolution and unconstrained viewpoints. To enable fine-grained semantic alignment, we incorporated UCA (Wu et al., 2024), which enhances UCF-Crime with 23,542 linguistic annotations spanning 110.7 hours of the original footage. We further expanded the training distribution with XD-Violence (Wu et al., 2020), contributing 4,754 untrimmed videos totaling 217 hours, and ActivityNet Captions (Krishna et al., 2017), which adds 849 hours of footage across 20,000 videos. This diverse training corpus ensures the model is exposed to a wide spectrum of visual environments, action durations, and linguistic complexities.
Training Setup. The cross-modal encoder was fine-tuned on an NVIDIA A100 GPU (40 GB VRAM). We implemented a cosine learning rate decay policy (Loshchilov and Hutter, 2017b), which gradually reduces the learning rate to optimize convergence. The initial learning rate ($lr_{\text{base}}$) was set to $8\times 10^{-6}$, selected based on successful configurations in similar CLIP fine-tuning tasks (Rasheed et al., 2023; Ma et al., 2022b). We employed the AdamW optimizer (Loshchilov and Hutter, 2017a) with a decoupled weight decay parameter of $1\times 10^{-3}$. The learning rate optimization follows the cosine annealing schedule defined as:
$\mathrm{lr}(t)=\frac{1}{2}\,\mathrm{lr}_{\mathrm{base}}\left(1+\cos\!\left(\frac{t-t_{\mathrm{warm\text{-}up}}}{T_{\mathrm{total}}-t_{\mathrm{warm\text{-}up}}}\,\pi\right)\right)$ (4)
where $T_{\text{total}}$ represents the total number of training steps and $t_{\text{warm-up}}$ denotes the number of warm-up steps. Given that model performance is highly sensitive to input granularity and batch composition, we evaluated multiple configurations during training, utilizing batch sizes ranging from 5 to 100 and frame sampling intervals of 5 to 40 equidistant frames per video clip.
4.2 Annotations and Interaction Traces
We provide the dataset containing the anonymized interaction traces from human annotators collected during our study of AI-assisted dense temporal video annotation. The dataset captures not only the final human segmentations, but also the full edit lifecycle, including model pre-annotations, the annotation guide, editing actions, and session-level timing information. This resource and the code of the adapted user interface are intended to support reproducible benchmarking of Human-in-the-Loop workflows in video annotation, enabling analyses of efficiency, revision behavior, and potential anchoring effects beyond what can be inferred from final labels alone. Table 1 provides a statistical summary of the collected data.
Table 1: Detailed statistics of the annotation dataset and interaction traces.
Metric Value
Annotation & Segments
Total Annotators 18
Total Annotation Sessions 180
Total Labeled Segments 699
Semantics
Mean Classes per Video $3.01\pm 0.95$
Abnormal Event Segments 323 (48.2%)
Normal Event Segments 347
4.3 Temporal Analysis: Efficiency Gains
The primary objective of the hybrid workflow is to reduce the cognitive load and manual effort required for dense video labeling. We tracked the precise task duration—from initial load to submission—for all 180 annotation sessions and analyzed the time-on-task distribution (Figure 5), which follows a Gamma-like distribution characteristic of cognitive tasks (Luce, 1991). Statistical tests, detailed in Table 2, confirm the independence of the two distributions ($p<0.005$).